# Stacked Intelligent Metasurfaces for Task-Oriented Semantic Communications

### v1

@@ IV-A Simulation Settings
In this section, we use the MNIST dataset to evaluate the performance of the proposed SIM-based DNN for implementing image recognition tasks in the wave domain. It consists of approximately 70,000 images of handwritten digits from 0 to 9 with $28\times 28$ grayscale pixels. The cardinality ratio of the training set to the testing set is 6:1, and 20% of the training set is allocated for validation purposes. Moreover, the semantic communication system operates at $28$ GHz, corresponding to a wavelength of $\lambda=10.7$ mm. The thickness of the SIM is set to $D=10\lambda$, while the spacing between adjacent meta-atoms is a wavelength, yielding $S_{a}=\lambda^{2}$. Additionally, the horizontal distance from the transmitting antenna to the input layer of the SIM is set to $d_{t,s}=d_{s}$. The distance between adjacent receiving antennas is set to $d_{r}=\lambda/2$. For the wireless link, the Rician factor is set to $K=3$ dB, and the path loss from the SIM to the receiver is modeled as $p=C_{0}d_{s,r}^{-\gamma}$, where $C_{0}=-35$ dB denotes the path loss at a reference distance of 1 m. $d_{s,r}$ denotes the distance between the SIM and the receiving array, and $\gamma=2.8$ is the path loss exponent in our simulations [19]. Additionally, the transmit power and the noise power are set to $p_{t}=40$ dBm and $\sigma^{2}_{v}=-104$ dBm, respectively.
In terms of the training process of SIM-based DNN, we consider 100 epochs, and the mini-batch size $B$ is set to $64$. Moreover, the learning rate is set to $\alpha=0.001$, which will be gradually decreased by multiplying $\alpha$ with a decay factor $\iota_{f}=0.8$, if the loss function value does not vary in five consecutive epochs. The decay coefficients of the first-order and the second-order moment estimates are $\beta_{1}=0.9$ and $\beta_{2}=0.999$, respectively. The first-order moment vector $\mathbf{m}_{t}$ and the second-order moment vector $\mathbf{v}_{t}$ are initialized to $\mathbf{0}$.
@@
Figure 2: Recognition accuracy using the SIM-based DNN.
@@ IV-B Performance Evaluation
In Fig. 2(a), we evaluate the effect of the number $L$ of layers on the recognition accuracy, where we consider $N=441$, $d_{s,r}=10$ m. For each input image, we retain only the $21\times 21$ pixels in the center. For simplicity, $M=8$ antennas are deployed at the receiver for recognizing digits $0\sim 7$. As observed in Fig. 2(a), the recognition accuracy using the SIM-based DNN increases with the number of metasurface layers, thanks to the enhanced inference capability of the multi-layer diffractive architecture of the SIM for achieving more accurate beam steering.
@@
Fig. 2(b) shows the recognition accuracy under different numbers $N$ of meta-atoms on each layer, where we set $L=7$ and maintain the other parameters the same as Fig. 2(a). For each input image, we preprocess it to match the size of the source encoding layer by using the bilinear interpolation method [16]. Note that the recognition accuracy improves as $N$ increases, since a larger number of learnable neurons (meta-atoms) would enhance the representational capability of the SIM. Additionally, increasing the number of meta-atoms also increases the aperture of the output metasurface layer, thus having a higher Rayleigh resolution to distinguish the signals at different receiving antennas. Specifically, the seven-layer SIM with $N=25$ meta-atoms on each layer achieves a low recognition accuracy of only $24\%$ on the test set. As $N$ increases to $121$ and $441$, the recognition accuracy on the testing set becomes $77\%$ and $84\%$, respectively.
@@
Fig. 2(c) illustrates the recognition accuracy versus the training epoch as the distance $d_{s,r}$ between the SIM and receiver increases, with $N=441$ and other parameters unchanged. It is observed from Fig. 2(c) that a shorter propagation distance would improve the recognition accuracy. This is due to a smaller value of $d_{s,r}$ would cause stronger channel gain compared to having a longer propagation distance. Hence, the noise effects would become increasingly negligible. Secondly, as the receiving array moves closer to the transmitter, the channels spanning from the SIM to different antennas would become more distinguishable. As a result, decreasing the propagation distance from $d_{s,r}=50$ m to $d_{s,r}=10$ m enhances the recognition accuracy from $72\%$ to $84\%$.
@@
Fig. 2(d) illustrates the recognition accuracy under different numbers of image classes where we set $d_{s,r}=10$ m. Specifically, we examine nine scenarios by gradually increasing both the number of image classes as well as the number of receiving antennas. Observe from Fig. 2(c) that the recognition accuracy improves as the number of image classes decreases. This is because recognizing fewer handwritten digits imposes a moderate demand on the inference capability of the SIM-based DNN. As a result, it becomes easier to focus the beams on the corresponding antenna. Additionally, a faster convergence trend is observed as $M$ decreases.
@@
Finally, to better illustrate the effectiveness of the proposed SIM-based DNN, Fig. 3 depicts the input image and the corresponding energy distribution at the receiving array, while the confusion matrix of utilizing the SIM-based DNN to recognize ten handwritten digits is shown in Fig. 4. It is demonstrated that the DNN is capable of focusing the information-bearing signals passing through the SIM onto the corresponding antenna.
@@ V Conclusion
In this letter, we investigated a SIM-aided image recognition task-oriented semantic communication. We trained a SIM in front of the transmitting antenna to steer the signals passing from the image source encoding layer towards the corresponding antenna. As a result, the receiver identified the image class by probing the magnitude of the received signal. Extensive simulation results verify the effectiveness of utilizing the SIM-based DNN for implementing image recognition task-oriented semantic communications, achieving more than $90\%$ recognition accuracy.

### v2

@@ IV-A Simulation Settings
In this section, we use the MNIST dataset to evaluate the performance of the proposed SIM-based EMNN for implementing image recognition tasks in the wave domain. It consists of approximately 70,000 images of handwritten digits from 0 to 9 with $28\times 28$ grayscale pixels. The cardinality ratio of the training set to the testing set is 6:1, and 20% of the training set is allocated for validation purposes. Note that the noise is added to the received signal during the training process to improve the system’s robustness. Moreover, the SemCom system operates at $28$ GHz, corresponding to a wavelength of $\lambda=10.7$ mm. The thickness of the SIM is set to $D=10\lambda$, while the spacing between adjacent meta-atoms is a wavelength, yielding $S_{a}=\lambda^{2}$. Additionally, the horizontal distance from the transmitting antenna to the input layer of the SIM is set to $d_{t,s}=d_{s}$. The distance between adjacent receiving antennas is set to $d_{r}=\lambda/2$. For the wireless link, the Rician factor is set to $K=3$ dB, and the path loss from the SIM to the receiver is modeled as $p=C_{0}d_{s,r}^{-\gamma}$, where $C_{0}=-35$ dB denotes the path loss at a reference distance of 1 m. $d_{s,r}$ denotes the distance between the SIM and the receiving array, and $\gamma=2.8$ is the path loss exponent in our simulations [26]. Additionally, the transmit power and the noise power are set to $p_{t}=40$ dBm and $\sigma^{2}_{v}=-104$ dBm, respectively.
In terms of the training process of SIM-based EMNN, we consider 100 epochs, and the mini-batch size $B$ is set to $64$. Moreover, the learning rate is set to $\alpha=0.001$, which will be gradually decreased by multiplying $\alpha$ with a decay factor $\iota_{f}=0.8$, if the loss function value does not vary in five consecutive epochs. The decay coefficients of the first-order and the second-order moment estimates are $\beta_{1}=0.9$ and $\beta_{2}=0.999$, respectively. The first-order moment vector $\mathbf{m}_{t}$ and the second-order moment vector $\mathbf{v}_{t}$ are initialized to $\mathbf{0}$.
@@
Figure 2: Recognition accuracy using the SIM-based EMNN.
@@ IV-B Performance Evaluation
In Fig. 2(a), we evaluate the effect of the number $L$ of layers on the recognition accuracy, where we consider $N=441$, $d_{s,r}=10$ m. For each input image, we retain only the $21\times 21$ pixels in the center. For simplicity, $M=8$ antennas are deployed at the receiver for recognizing digits $0\sim 7$. As observed in Fig. 2(a), the recognition accuracy using the SIM-based EMNN increases with the number of metasurface layers, thanks to the enhanced inference capability of the multi-layer architecture for achieving more accurate beam steering.
@@
Fig. 2(b) shows the recognition accuracy under different numbers $N$ of meta-atoms on each layer, where we set $L=7$ and maintain the other parameters the same as Fig. 2(a). For each input image, we preprocess it to match the size of the source encoding layer by using the bilinear interpolation method [22]. Note that the recognition accuracy improves as $N$ increases, since a larger number of learnable neurons (meta-atoms) would enhance the representational capability of the SIM. Additionally, increasing the number of meta-atoms also increases the aperture of the output metasurface layer, thus having a higher Rayleigh resolution to distinguish the signals at different receiving antennas. Specifically, the seven-layer SIM with $N=25$ meta-atoms on each layer achieves a low recognition accuracy of only $32\%$ on the testing set. As $N$ increases to $121$ and $441$, the recognition accuracy on the testing set becomes $77\%$ and $84\%$, respectively.
According to the simulation results in Fig. 2(a)-(b), increasing the value of $L$ and $N$ enhances recognition accuracy. However, the improvement of recognition accuracy is relatively small when $L$ exceeds $7$, as shown in Fig. 2(a). Similarly, Fig. 2(b) demonstrates the performance plateaus when $N$ exceeds 441. These diminishing returns suggest choosing proper SIM parameters for achieving the desired performance with minimum hardware cost.
@@
Fig. 2(c) illustrates the recognition accuracy versus the training epoch as the distance $d_{s,r}$ between the SIM and receiver increases, with $N=441$ and other parameters unchanged. It is observed from Fig. 2(c) that a shorter propagation distance would improve the recognition accuracy. This is due to a smaller value of $d_{s,r}$ would cause stronger channel gain compared to that with a longer propagation distance. Hence, the noise effects would become increasingly negligible. Secondly, as the receiving array moves closer to the transmitter, the channels from the SIM to different antennas would become more distinguishable. As a result, decreasing the propagation distance from $d_{s,r}=50$ m to $d_{s,r}=10$ m enhances the recognition accuracy from $72\%$ to $84\%$.
@@
Fig. 2(d) illustrates the recognition accuracy under different numbers of image classes where we set $d_{s,r}=10$ m. Specifically, we examine nine scenarios by gradually increasing both the number of image classes as well as the number of receiving antennas. Note that recognizing more handwritten digits requires the SIM to accurately focus multiple beams on an increased number of receiving antennas. As observed from Fig. 2(d), the recognition accuracy becomes worse as the number of image classes increases. Additionally, a faster convergence trend is observed as $M$ decreases.
@@
Finally, to better illustrate the effectiveness of the proposed SIM-based EMNN, Fig. 3 depicts the input image and the corresponding energy distribution at the receiving array, while the confusion matrix of utilizing the SIM-based EMNN to recognize ten handwritten digits is shown in Fig. 4. It is demonstrated that the EMNN is capable of focusing the information-bearing signals passing through the SIM onto the corresponding antenna.
@@ V Conclusion
In this letter, we investigated a SIM-aided image recognition task-oriented SemCom. We trained a SIM in front of the transmitting antenna to steer the signals passing from the image source encoding layer towards the corresponding antenna. As a result, the receiver identified the image class by probing the magnitude of the received signal. Extensive simulation results verify the effectiveness of the SIM-based EMNN in achieving over 90% recognition accuracy for image recognition task-oriented SemComs. Combined with traditional modulated signals, SIM is expected to further achieve multimodal semantic information transmission.